40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Dec 16 18:34:40 2015, gautam, Update, Green Locking, Green beat channels temporarily set up as IR beat channels 
    Reply  Wed Dec 16 23:15:28 2015, ericq, Update, Green Locking, Green beat channels temporarily set up as IR beat channels 2015-12-16_AUXfreerunning.pdf
Message ID: 11888     Entry time: Wed Dec 16 23:15:28 2015     In reply to: 11887
Author: ericq 
Type: Update 
Category: Green Locking 
Subject: Green beat channels temporarily set up as IR beat channels 

With the IR beats going to the nominal ALS channels as Gautam left them, we're able to measure the free running frequency noise of the end AUX lasers. 

Specifically, the end shutters are closed, leaving the AUX lasers free running. The IR beats then consist of this free running light beating with the PSL light, and the ALS phase trackers give a calibrated frequency noise spectrum. I've stabilized the PSL light by locking the laser to the Y arm via MC2 acutation, so the free running AUX laser noise should dominate by a lot above the suspension resonances. This also has the benefit of giving me the use of the CAL'd Y arm displacement as a sanity check. 

At this point in time, it looks like the X laser is close to 10x noisier than the Y laser, though it does seem to be at the rule-of-thumb "10kHz/rtHz at 100Hz" level. 

Attachment 1: 2015-12-16_AUXfreerunning.pdf  81 kB  | Hide | Hide all
2015-12-16_AUXfreerunning.pdf
ELOG V3.1.3-